OpenStudy (anonymous):

Hi again. I need help with this problem. Sorry if I'm bugging anyone with these problems. please help. Solve for x: 4(x-5)= x+8 Please show your steps. I need to know how you did it. I have that problem wrong before and I need to know my mistake.

OpenStudy (anonymous):

just do simply

OpenStudy (anonymous):

?

jhonyy9 (jhonyy9):

4x-20=x+8 3x=28 x=28/3

OpenStudy (anonymous):

try to collect variable in one side of = so here 4x-20= x+8 4x-x=8+20 3x= 28 so x= 28/3

OpenStudy (anonymous):

Ok thanks. :)

OpenStudy (anonymous):

Thats not the final answer. its 9 1/3